# Higher derivative heterotic supergravity on a torus and supersymmetry

### v2

@@ 2 Torus reduction of heterotic supergravity
$\Omega_{\pm}=\omega\pm\frac{1}{2}\tilde{\mathcal{H}},\qquad\tilde{\mathcal{H}}^{AB}\equiv\tilde{H}_{M}{}^{AB}dx^{M},$ (2.3)
and $\tilde{H}_{MNP}$ is the three-form with Chern-Simons addition
@@
Note that $\tilde{H}$ satisfies the Bianchi identity
@@ 2.3 The two-derivative reduction and supersymmetry
$\mathcal{E}_{F^{+},\alpha a}=$ $\;e^{2\varphi}\,\nabla_{\beta}(e^{-2\varphi}F_{\alpha\beta}^{(+)\,a})-Q_{\beta}^{(++)\,ab}F_{\alpha\beta}^{(+)\,b}-\dfrac{1}{2}h_{\alpha\beta\delta}+P_{\beta}^{(+-)\,ac}F_{\alpha\beta}^{(-)\,c}$
$\mathcal{E}_{F^{-},\alpha a}=$ $\;e^{2\varphi}\,\nabla_{\beta}(e^{-2\varphi}F_{\alpha\beta}^{(-)\,a})-Q_{\beta}^{(--)\,ab}F_{\alpha\beta}^{(-)\,b}+\dfrac{1}{2}h_{\alpha\beta\delta}+P_{\beta}^{(+-)\,ca}F_{\alpha\beta}^{(+)\,c}$
@@
$\mathcal{E}_{h,\alpha\beta}=$ $\,e^{2\phi}\nabla_{\gamma}(e^{-2\phi}h_{\alpha\beta\gamma})$
ϕ=subscriptitalic-$\mathcal{E}_{\phi}=$ $\;R(\omega)+4\Box\varphi-4(\nabla_{\alpha}\varphi)^{2}-\dfrac{1}{8}\left(F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\alpha\beta}+F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\alpha\beta}\right)-\dfrac{1}{12}h_{\alpha\beta\gamma}^{2}$
@@ 4.1 The Lagrangian with a torsion-free connection
$+\dfrac{3}{8}h_{\alpha\beta\epsilon}h_{\gamma\delta\epsilon}F_{\alpha\gamma}^{(-)\,a}F_{\beta\delta}^{(-)\,a}-\dfrac{3}{8}h_{\alpha\beta\epsilon}h_{\gamma\delta\epsilon}F_{\alpha\gamma}^{(-)\,a}F_{\beta\delta}^{(-)\,a},$ (4.13)
@@ 5 Discussion
Clearly, the combination of T-duality invariance and world-sheet parity is more restrictive than T-duality alone. Since the type II strings are invariant under world-sheet parity (along with space-time parity in the IIB case), it would be interesting to see if this combination of symmetries could shed further light on the structure of T-duality invariant couplings of the type II string from an $O(d)_{-}\times O(d)_{+}$ point of view. A full consideration would involve the Ramond-Ramond sector as well, in which case the $O(d)_{-}\times O(d)_{+}$ symmetry would be enhanced to the bottom of the U-duality group. A natural way to pursue this investigation could be through the use of exceptional field theory Hohm:2013pua ; Hohm:2013vpa ; Hohm:2013uia ; Hohm:2014fxa .

### v3

@@ 2 Torus reduction of heterotic supergravity
$\Omega_{\pm}=\omega\pm\frac{1}{2}\tilde{\mathcal{H}},\qquad\tilde{\mathcal{H}}^{AB}\equiv\tilde{H}_{M}{}^{AB}dx^{M}.$ (2.3)
In the above, $\tilde{H}_{MNP}$ is the three-form with Chern-Simons addition
@@
In particular, $\tilde{H}$ satisfies the Bianchi identity
@@
Note that $\tilde{H}$ appears on both sides of (2.4) since it enters implicitly on the right-hand side through the definition of the torsionful connection $\Omega_{+}$. This structure can be seen from the supersymmetric construction of Bergshoeff:1988nn ; Bergshoeff:1989de as well as from $\sigma$-model anomaly considerations Sen:1985tq . The proper way to make sense of this expression for $\tilde{H}$ is as an iterative expansion in $\alpha^{\prime}$. Since we work only to first order in $\alpha^{\prime}$, it is sufficient for our purposes to take the leading order expression for the torsionful connection, so that
$\tilde{H}=H-\frac{\alpha^{\prime}}{4}\omega_{3L}(\omega+{\textstyle\frac{1}{2}}\mathcal{H}),\qquad\mathcal{H}=H_{M}{}^{AB}dx^{M}.$ (2.7)
Similar considerations apply when working with the torsionful Riemann tensor.
@@ 2.3 The two-derivative reduction and supersymmetry
$\mathcal{E}_{F^{+},\alpha a}=$ $\;e^{2\varphi}\,\nabla_{\beta}(e^{-2\varphi}F_{\alpha\beta}^{(+)\,a})-Q_{\beta}^{(++)\,ab}F_{\alpha\beta}^{(+)\,b}-\dfrac{1}{2}h_{\alpha\beta\delta}F_{\beta\delta}^{(+)\,a}+P_{\beta}^{(+-)\,ac}F_{\alpha\beta}^{(-)\,c}$
$\mathcal{E}_{F^{-},\alpha a}=$ $\;e^{2\varphi}\,\nabla_{\beta}(e^{-2\varphi}F_{\alpha\beta}^{(-)\,a})-Q_{\beta}^{(--)\,ab}F_{\alpha\beta}^{(-)\,b}+\dfrac{1}{2}h_{\alpha\beta\delta}F_{\beta\delta}^{(-)\,a}+P_{\beta}^{(+-)\,ca}F_{\alpha\beta}^{(+)\,c}$
@@
$\mathcal{E}_{h,\alpha\beta}\;\;\,=$ $\;e^{2\varphi}\nabla_{\gamma}(e^{-2\varphi}h_{\alpha\beta\gamma})$
$\mathcal{E}_{\phi}\ \ \ \ =$ $\;R(\omega)+4\Box\varphi-4(\nabla_{\alpha}\varphi)^{2}-\dfrac{1}{8}\left(F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\alpha\beta}+F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\alpha\beta}\right)-\dfrac{1}{12}h_{\alpha\beta\gamma}^{2}$
@@ 4.1 The Lagrangian with a torsion-free connection
$+\dfrac{3}{8}h_{\alpha\beta\epsilon}h_{\gamma\delta\epsilon}F_{\alpha\gamma}^{(-)\,a}F_{\beta\delta}^{(-)\,a}-\dfrac{3}{8}h_{\alpha\beta\epsilon}h_{\gamma\delta\epsilon}F_{\alpha\gamma}^{(+)\,a}F_{\beta\delta}^{(+)\,a},$ (4.13)
@@ 5 Discussion
Clearly, the combination of T-duality invariance and world-sheet parity is more restrictive than T-duality alone. Since the type II strings are invariant under world-sheet parity (along with space-time parity in the IIB case), it would be interesting to see if this combination of symmetries could shed further light on the structure of T-duality invariant couplings of the type II string from an $O(d)_{-}\times O(d)_{+}$ point of view. A full consideration would involve the Ramond-Ramond sector as well, in which case the $O(d)_{-}\times O(d)_{+}$ symmetry would be enhanced to the bottom of the U-duality group. A natural way to pursue this investigation could be through the use of exceptional field theory Hohm:2013pua ; Hohm:2013vpa ; Hohm:2013uia ; Hohm:2014fxa . Note that U-duality also arises in bosonic and heterotic compactifications to three dimensions, and the interplay between $\mathcal{O}(\alpha^{\prime})$ corrections and U-duality has been investigated in Eloy:2022vsq for this case.